# Shifted Interpolation for Differential Privacy

### v1

@@ C.1 Shifted interpolation for contractive noisy iterations
[…]
@@
[…]
@@
In the setting of Theorem C.3, additionally assume that $10<c<$ 1. Then
@@
In Theorem C.3, we can take $\tau=0,z_{\tau}=0$ and $s_{k+1}\equiv s$. Then the values of $\{\lambda_{k}\},\{z_{k}\},\{a_{k}\}$ obtained from the elementary optimization problem (C.6) yield the desired result. Finally, for the equality case, by direct calculation we have $X_{t}\sim\mathcal{N}(0,\frac{1-c^{2t}}{1-c^{2}}\sigma^{2}I_{d})$ and $X^{\prime}_{t}\sim\mathcal{N}(\frac{1-c^{t}}{1-c}sv,\frac{1-c^{2t}}{1-c^{2}}\sigma^{2}I_{d})$, giving
@@
This follows as a direct corollary of Theorem C.5 with $\phi_{k}(x)\equiv\phi(x)=x-\frac{\eta}{n}\sum_{i=1}^{n}\nabla f_{i}(x)$ and $\phi^{\prime}_{k}(x^{\prime})\equiv\phi^{\prime}(x^{\prime})=x^{\prime}-\frac{\eta}{n}\sum_{i=1}^{n}\nabla f^{\prime}_{i}(x^{\prime})$. For this application, consider parameters $c=\max\{|1-\eta m|,|1-\eta M|\}<1$ (by 2.8), $s\leftarrow\eta L/n$, and $\sigma\leftarrow\eta\sigma$ (rescaling to simplify notation). The equality case is a straightforward calculation in the setting that $\mathcal{K}=\mathbb{R}^{d}$, $X_{0}=0$, $\nabla f_{i}(x)=mx$ for all $i\in[n]$, and $\nabla f^{\prime}_{i}(x)$ defined as $mx$ for $i\neq i^{*}$, and otherwise $mx+Lv$ for some unit vector $v$. ∎
@@
In the setting of Theorem C.3, additionally assume that $\mathcal{K}$ has a finite diameter $D$. Then for any integer $0\leq\tau<t$,
@@
From $\lVert{X_{\tau}-\widetilde{X}_{\tau}}\rVert\leq D$ for all $\tau$, in Theorem C.3 we can take $z_{\tau}=D,s_{k}\equiv s$, and $c=1$. The values of $\{\lambda_{k}\},\{z_{k}\},\{a_{k}\}$ are obtained by analyzing the following elementary optimization problem (C.8). ∎
@@
[…]
@@
where the first inequality holds from the post-processing inequality with $\phi_{k}\equiv\phi^{\prime}_{k}$ for all $k=t+j^{*}-l+1,\dots,t$, and the second inequality holds from $\lVert{\phi_{t+j^{*}-l}-\phi^{\prime}_{t+j^{*}-l}}\rVert_{\infty}\leq\frac{\eta L}{b}$ with 3.1.
In general, $\phi_{k+1}=\phi^{\prime}_{k+1}$ when $r=r(k)=k+1-l\lfloor\frac{k}{l}\rfloor$ is not equal to $j^{*}$; otherwise $\lVert{\phi_{k+1}-\phi^{\prime}_{k+1}}\rVert_{\infty}\leq\frac{\eta L}{b}$. Thus, in Theorem C.3 we can take $\tau=0,z_{\tau}=0$ and $s_{k+1}=s\bm{1}_{\{r=j^{*}\}}$ where $s=\frac{\eta L}{b}$. Using $\{\lambda_{k}\},\{z_{k}\},\{a_{k}\}$ obtained from the following result (C.10),
@@ C.4.2 Proofs of new results
[…]
@@
Figure 5: Illustration of shifted interpolated processes in the proofs of Theorem 4.9 (left) and Theorem 4.10 (right). The dotted lines denote the alternative updates based on $V_{k}=0$ or $1$; each interpolated process uses the same value of $V_{k}$ as expressed in the figure. In Theorem 4.9, we build two processes, each of which tracks its corresponding original process. In Theorem 4.10, only one process is built and it inherits the identical deviation based on the realization of $V_{k}$.

### v2

@@ C.1 Shifted interpolation for contractive noisy iterations
Consider processes $\{X_{k}\}$ and $\{X_{k}^{\prime}\}$ corresponding respectively to $\text{CNI}(X_{0},\{\phi_{k}\}_{k\in[t]},\{\xi_{k}\}_{k\in[t]},\mathcal{K})$ and $\text{CNI}(X_{0}^{\prime},\{\phi_{k}^{\prime}\}_{k\in[t]},\{\xi_{k}\}_{k\in[t]},\mathcal{K})$. The shifted interpolated process between these two CNI is the auxiliary process $\{\widetilde{X}_{k}\}$ satisfying $\widetilde{X}_{\tau}=X^{\prime}_{\tau}$ and
@@
Let $X_{t}$ and $X^{\prime}_{t}$ respectively be the output of $\text{CNI}(X_{0},\{\phi_{k}\}_{k\in[t]},\{\mathcal{N}(0,\sigma^{2}I_{d})\}_{k\in[t]},\mathcal{K})$ and $\text{CNI}(X_{0},\{\phi^{\prime}_{k}\}_{k\in[t]},\{\mathcal{N}(0,\sigma^{2}I_{d})\}_{k\in[t]},\mathcal{K})$ such that each $\phi_{k},\phi^{\prime}_{k}$ is $c$-Lipschitz and $\lVert{\phi_{k}(x)-\phi^{\prime}_{k}(x)}\rVert\leq s_{k}$ for all $x$ and $k\in[t]$. Then for any intermediate time $\tau$ and shift parameters $\lambda_{\tau+1},\dots,\lambda_{t}\in[0,1]$ with $\lambda_{t}=1$,
@@
In the setting of Theorem C.3, additionally assume that $10<c<$ 1 and $s_{k}\equiv s$. Then
@@
In Theorem C.3, we can take $\tau=0$ and $z_{\tau}=0$. Then the values of $\{\lambda_{k}\},\{z_{k}\},\{a_{k}\}$ obtained from the elementary optimization problem (C.6) yield the desired result. Finally, for the equality case, by direct calculation we have $X_{t}\sim\mathcal{N}(0,\frac{1-c^{2t}}{1-c^{2}}\sigma^{2}I_{d})$ and $X^{\prime}_{t}\sim\mathcal{N}(\frac{1-c^{t}}{1-c}sv,\frac{1-c^{2t}}{1-c^{2}}\sigma^{2}I_{d})$, giving
@@
This follows as a direct corollary of Theorem C.5 with $\phi_{k}(x)\equiv\phi(x)=x-\frac{\eta}{n}\sum_{i=1}^{n}\nabla f_{i}(x)$ and $\phi^{\prime}_{k}(x^{\prime})\equiv\phi^{\prime}(x^{\prime})=x^{\prime}-\frac{\eta}{n}\sum_{i=1}^{n}\nabla f^{\prime}_{i}(x^{\prime})$. For this application, consider parameters $c=\max\{|1-\eta m|,|1-\eta M|\}<1$ (by 2.8), $s\leftarrow\eta L/n$, and $\sigma\leftarrow\eta\sigma$ (rescaling to simplify notation). The equality case is a straightforward calculation in the setting that $\mathcal{K}=\mathbb{R}^{d}$, $X_{0}=0$, $\nabla f_{i}(x)=mx$ for all $i\in[n]$, and $\nabla f^{\prime}_{i}(x)$ defined as $mx$ for $i\neq i^{*}$, and otherwise $mx-Lv$ for some unit vector $v$. ∎
@@
In the setting of Theorem C.3, additionally assume that $\mathcal{K}$ has a finite diameter $D$ and $s_{k}\equiv s$. Then for any integer $0\leq\tau<t$,
@@
From $\lVert{X_{\tau}-\widetilde{X}_{\tau}}\rVert\leq D$ for all $\tau$, in Theorem C.3 we can take $z_{\tau}=D$ and $c=1$. The values of $\{\lambda_{k}\},\{z_{k}\},\{a_{k}\}$ are obtained by analyzing the following elementary optimization problem (C.8). ∎
@@
For $t=lE$ and $k=0,1,\dots,t-1$, let $B_{1},\dots,B_{l}$ be a fixed partition of $[n]$ with size $b$, and define $\phi_{k+1}(x)=x-\frac{\eta}{b}\sum_{i\in B_{r}}\nabla f_{i}(x)$ and $\phi^{\prime}_{k+1}(x^{\prime})=x^{\prime}-\frac{\eta}{b}\sum_{i\in B_{r}}\nabla f^{\prime}_{i}(x^{\prime})$ where $r=k+1-l\lfloor\frac{k}{l}\rfloor$. Then the $f$-DP of NoisyCGD is equal to that between $X_{t}=\text{CNI}(X_{0},\{\phi_{k}\}_{k\in[t]},\{\mathcal{N}(0,\eta^{2}\sigma^{2}I_{d})\}_{k\in[t]},\mathcal{K})$ and $X^{\prime}_{t}=\text{CNI}(X_{0},\{\phi^{\prime}_{k}\}_{k\in[t]},\{\mathcal{N}(0,\eta^{2}\sigma^{2}I_{d})\}_{k\in[t]},\mathcal{K})$.
@@
where the first inequality holds from the post-processing inequality with $\phi_{k}\equiv\phi^{\prime}_{k}$ for all $k=t+j^{*}-l+1,\dots,t$, and the second inequality holds from $\lVert{\phi_{t+j^{*}-l}(x)-\phi^{\prime}_{t+j^{*}-l}(x)}\rVert\leq\frac{\eta L}{b}$ for all $x$ with 3.1.
In general, $\phi_{k+1}=\phi^{\prime}_{k+1}$ when $r=r(k)=k+1-l\lfloor\frac{k}{l}\rfloor$ is not equal to $j^{*}$; otherwise $\lVert{\phi_{k+1}(x)-\phi^{\prime}_{k+1}(x)}\rVert\leq\frac{\eta L}{b}$ for all $x$. Thus, in Theorem C.3 we can take $\tau=0,z_{\tau}=0$ and $s_{k+1}=s\bm{1}_{\{r=j^{*}\}}$ where $s=\frac{\eta L}{b}$. Using $\{\lambda_{k}\},\{z_{k}\},\{a_{k}\}$ obtained from the following result (C.10),
@@ C.4.2 Proofs of new results
$R(s,\sigma,p)=\inf\{T(VW+Z,VW^{\prime}+Z):$ $\ V\sim\text{Ber}(p),\lVert{W}\rVert,\lVert{W^{\prime}}\rVert\leq s,Z\sim\mathcal{N}(0,\sigma^{2}I_{d})\}\,$
@@
Figure 5: Illustration of shifted interpolated processes in the proofs of Theorem 4.9 (left) and Theorem 4.10 (right). The solid lines denote the updates based on the realized values of $\{V_{k}\}$, and the dashed lines denote the alternative updates based on their unrealized values; each interpolated process uses the same (coupled) values of $\{V_{k}\}$ as expressed in the figure. In Theorem 4.9, we build two processes, each of which tracks its corresponding original process. In Theorem 4.10, only one process is built and it inherits the identical deviation based on the realizations of $\{V_{k}\}$.
@@
C.4.3 Choice of $t-\tau$ based on approximation
Since Theorem 4.9 and Theorem 4.10 hold for every $t-\tau$, we can calculate the corresponding $f$-DP bound for each $t-\tau$ and then take the pointwise maximum as a valid privacy guarantee. However, this may be computationally burdensome if $t$ is large. One way to bypass this calculation is to approximate the composition of subsampled Gaussian mechanisms via CLT (A.5), where the resulting $f$-DP bound becomes a GDP bound and thus optimization over $t-\tau$ is analytically tractable.
Proposition C.14.
In the setting of Theorem 4.9, by choosing (modulo floor or ceiling)
$t-\tau=-\frac{\log\frac{b^{2}\sigma(1-c)}{2\sqrt{2}nL\sqrt{\log(1/c)}}\sqrt{e^{4L^{2}/(b\sigma)^{2}}\Phi(\frac{3L}{b\sigma})+3\Phi(-\frac{L}{b\sigma})-2}}{\log(1/c)}-1$
NoisySGD is approximately $\mu$-GDP, where
$\mu=\sqrt{8\left(\frac{L}{b\sigma}\frac{c^{t-\tau+1}}{1-c}\right)^{2}+\frac{2b^{2}}{n^{2}}(t-\tau)(e^{4L^{2}/(b\sigma)^{2}}\Phi(\frac{3L}{b\sigma})+3\Phi(-\frac{L}{b\sigma})-2)}\,.$ (C.6)
Proof.
By A.5,
$C_{b/n}(G(\frac{2\sqrt{2}L}{b\sigma}))\otimes C_{b/n}(G(\frac{2L}{b\sigma}))^{\otimes(t-\tau)}\approx G\left(\sqrt{2}\frac{b}{n}\sqrt{(t-\tau)(e^{4L^{2}/(b\sigma)^{2}}\Phi(\frac{3L}{b\sigma})+3\Phi(-\frac{L}{b\sigma})-2)}\right)\,.$
Also, by bounding
$\frac{2\sqrt{2}L}{b\sigma}\frac{c^{t-\tau+1}-c^{t}}{1-c}\leq\frac{2\sqrt{2}L}{b\sigma}\frac{c^{t-\tau+1}}{1-c}$
we obtain an approximate lower bound $G(\mu)$ of the form (C.6). As a function of $t-\tau\in(0,\infty)$ it is convex, and the first-order optimality condition provides the stated formula for $t-\tau$. ∎
Proposition C.15.
In the setting of Theorem 4.10, by choosing (modulo floor or ceiling)
$t-\tau=\frac{Dn}{b\eta\sigma\sqrt{e^{8L^{2}/(b\sigma)^{2}}\Phi(\frac{3\sqrt{2}L}{b\sigma})+3\Phi(-\frac{\sqrt{2}L}{b\sigma})-2}}$
NoisySGD is approximately $\mu$-GDP, where
$\mu=\sqrt{\frac{2D^{2}}{\eta^{2}\sigma^{2}(t-\tau)}+2\frac{b^{2}}{n^{2}}(t-\tau)(e^{8L^{2}/(b\sigma)^{2}}\Phi(\frac{3\sqrt{2}L}{b\sigma})+3\Phi(-\frac{\sqrt{2}L}{b\sigma})-2)}\,.$ (C.7)
Proof.
As in the proof of C.14, the CLT approximation of $C_{b/n}(G(\frac{2\sqrt{2}L}{b\sigma}))^{\otimes(t-\tau)}$ provides a lower bound $G(\mu)$ of the form (C.7), which is a convex function in $t-\tau$; the first-order optimality condition yields the stated result. ∎
C.5 Lower bounds
Here we elaborate on lower bounds for the privacy loss (i.e., upper bounds on the $f$-DP guarantee) that complement our results in §4. Note that an exactly matching bound for NoisyGD in the strongly convex setting was obtained in Theorem 4.2, and an asymptotically matching bound for NoisySGD in the constrained convex setting was obtained in [AT22, ABT24]. Below, we present results for the other related settings using similar techniques. For the strongly convex setting, these lower bounds are built based on convex quadratics which yield iterates with explicit Gaussians; and for the constrained convex setting, these are obtained by comparing symmetric and biased (projected) Gaussians. We refer the readers to [AT22, ABT24] for further discussion about these constructions.
Theorem C.16.
Consider the setting of Theorem 4.3 or Theorem 4.6. There exist universal constants $00<c_{0}<1/5,c_{1}>$ 0 such that if $\sigma^{2}\leq c_{0}\frac{LD}{\eta n}$ and $\mu=c_{1}\frac{1}{\sigma}\sqrt{\frac{LD}{\eta n}}$, then
  • (a)
    NoisyGD is not $\mu$-GDP for all $t\geq\frac{Dn}{\eta L}\geq\frac{1}{2}$.
  • (b)
    NoisyCGD is not $\mu$-GDP for all $E\geq\frac{Db}{\eta L}\geq\frac{1}{2}$.
Proof.
  • (a)
    For NoisyGD, let $\mu_{0}=\tfrac{\mu}{c_{1}}=\tfrac{1}{\sigma}\sqrt{\tfrac{LD}{\eta n}}$. Consider $d=1$ and loss functions such that $\nabla f_{i}(x)=0$ for all $i\in[n]$, $\nabla f^{\prime}_{i}(x)=0$ for all $i\neq i^{*}$ and $\nabla f^{\prime}_{i^{*}}(x)=-L$ [^12] For general $d>1$, a similar argument (with slightly different constants) can be made by considering $\nabla f^{\prime}_{i^{*}}(x)=-Le_{1}$ and $\mathcal{K}=[-\Theta(D),\Theta(D)]\times[-\Theta(D/\sqrt{d-1}),\Theta(D/\sqrt{d-1})]^{d-1}$ (constant factors chosen such that $\mathcal{K}$ has diameter $D$). Also, let $X_{0}=0$ and $\mathcal{K}=[-\tfrac{D}{2},\tfrac{D}{2}]$. Note that by A.3, a $\mu$-GDP algorithm is $(\mu^{2},\Phi(-\tfrac{\mu}{2}))$-DP. We will show that for $E=[-\tfrac{D}{2},0]$,
    $\mathbb{P}(X_{t}\in E)$ $=\frac{1}{2}$
    $\mathbb{P}(X^{\prime}_{t}\in E)$ $<\exp(-\mu^{2})(\frac{1}{2}-\Phi(-\frac{\mu}{2}))$
    which implies that NoisyGD is not $(\mu^{2},\Phi(-\tfrac{\mu}{2}))$-DP and thus NoisyGD is not $\mu$-GDP. First, recall that
    $X_{k+1}$ $=\Pi_{\mathcal{K}}(X_{k}+Z_{k+1})$
    $X^{\prime}_{k+1}$ $=\Pi_{\mathcal{K}}(X^{\prime}_{k}+\frac{\eta L}{n}+Z^{\prime}_{k+1})$
    where $Z_{k+1},Z^{\prime}_{k+1}\sim\mathcal{N}(0,\eta^{2}\sigma^{2})$. Since the distribution of $X_{k}$ is symmetric for all $k$, $\mathbb{P}(X_{t}\in E)=\tfrac{1}{2}$. On the other hand, for $t_{0}=t-\lceil 0.8\tfrac{Dn}{\eta L}\rceil+1$ consider a process $\{X^{\prime\prime}_{k}\}_{t_{0}\leq k\leq t}$ such that $X^{\prime\prime}_{t_{0}}=-\tfrac{D}{2}$ and
    $X^{\prime\prime}_{k+1}=\min\{X^{\prime\prime}_{k}+\frac{\eta L}{n}+Z^{\prime}_{k+1},\frac{D}{2}\}\,.$
    Then inductively, $\mathbb{P}(X^{\prime}_{k}\leq z)\leq\mathbb{P}(X^{\prime\prime}_{k}\leq z)$ for all $z$. Letting $E_{0}=\{\max_{t_{0}\leq k\leq t}\sum_{j=t_{0}}^{k}Z_{j}\leq 0.1D\}$, by Doob’s submartingale inequality we have
    $\mathbb{P}(E_{0}^{c})\leq\exp(-\frac{(0.1D)^{2}}{2\times\lceil 0.8\frac{Dn}{\eta L}\rceil\times(\eta\sigma)^{2}})\leq\exp(-\frac{0.01LD}{5.6n\eta\sigma^{2}})=\exp(-\frac{0.01}{5.6}\mu_{0}^{2})\,.$
    Also, conditioning on $E_{0}$, $X^{\prime\prime}_{t}=-\tfrac{D}{2}+\tfrac{\eta L}{n}\times\lceil 0.8\tfrac{Dn}{\eta L}\rceil+\sum_{j=t_{0}}^{t}Z_{j}\geq 0.3D+\sum_{j=t_{0}}^{t}Z_{j}$. Thus
    $\mathbb{P}(X^{\prime}_{t}\notin E)$ $\geq\mathbb{P}(X^{\prime\prime}_{t}>0)$
    ($\geq\mathbb{P}(\{X^{\prime\prime}_{t}>0\}\cap E_{0})$
    ($\geq\mathbb{P}(\{0.3D+\sum_{j=t_{0}}^{t}Z_{j}>0\}\cap E_{0})$
    $\geq\mathbb{P}(0.3D+\sum_{j=t_{0}}^{t}Z_{j}>0)-\exp(-\frac{0.01}{5.6}\mu_{0}^{2})$
    $\geq\Phi(\frac{0.3}{\sqrt{2.8}}\mu_{0})-\exp(-\frac{0.01}{5.6}\mu_{0}^{2})$
    $\geq 1-\exp(-\frac{0.9}{5.6}\mu_{0}^{2})-\exp(-\frac{0.01}{5.6}\mu_{0}^{2})$
    where the penultimate inequality is from that $\sum_{j=t_{0}}^{t}Z_{j}$ is a mean zero Gaussian with variance $\lceil 0.8\tfrac{Dn}{\eta L}\rceil\eta^{2}\sigma^{2}\leq\tfrac{2.8Dn\eta\sigma^{2}}{L}=\tfrac{2.8D^{2}}{\mu_{0}^{2}}$, and the last inequality is from $\Phi(x)\geq 1-\exp(-\tfrac{1}{2}x^{2})$ for all $x\geq\tfrac{1}{\sqrt{2\pi}}$ (with $0.3\mu_{0}/\sqrt{2.8}\geq 0.3/\sqrt{2.8c_{0}}\geq 1/\sqrt{2\pi}$). By taking sufficiently small $c_{1}<\sqrt{\tfrac{0.01}{5.6}}$ and $c_{0}<c_{1}^{2}$ such that
    $\exp(-\frac{0.9}{5.6}\mu_{0}^{2})+\exp(-\frac{0.01}{5.6}\mu_{0}^{2})\leq\exp(-c_{1}^{2}\mu_{0}^{2})(\frac{1}{2}-\Phi(-\frac{1}{2}))$
    for all $\mu_{0}^{2}\geq 1/c_{0}$, we have
    $\exp(-\mu^{2})(\frac{1}{2}-\Phi(-\frac{\mu}{2}))$ $=\exp(-c_{1}^{2}\mu_{0}^{2})(\frac{1}{2}-\Phi(-\frac{c_{1}\mu_{0}}{2}))$
    $\geq\exp(-c_{1}^{2}\mu_{0}^{2})(\frac{1}{2}-\Phi(-\frac{c_{1}}{2\sqrt{c_{0}}}))$
    $>\exp(-c_{1}^{2}\mu_{0}^{2})(\frac{1}{2}-\Phi(-\frac{1}{2}))$
    $\geq\exp(-\frac{0.9}{5.6}\mu_{0}^{2})+\exp(-\frac{0.01}{5.6}\mu_{0}^{2})$
    $\geq\mathbb{P}(X^{\prime}_{t}\in E)$
    as desired.
  • (b)
    The proof for NoisyCGD is similar to that for NoisyGD (recall that $t=lE$ and $n=lb$); consider the same loss functions, initialization, constraint set with $i^{*}\in B_{l}$. Then
    $X_{k+1}$ $=\Pi_{\mathcal{K}}(X_{k}+Z_{k+1})$
    $X^{\prime}_{k+1}$ $=\Pi_{\mathcal{K}}(X^{\prime}_{k}+\frac{\eta L}{b}\bm{1}_{\{r(k)=l\}}+Z^{\prime}_{k+1})$
    where $r(k)=k+1-l\lfloor\tfrac{k}{l}\rfloor$. For $t_{0}=l(E-\lceil 0.8\tfrac{Db}{\eta L}\rceil)+1$, consider a process $\{X^{\prime\prime}_{k}\}_{t_{0}\leq k\leq t}$ such that $X^{\prime\prime}_{t_{0}}=-\tfrac{D}{2}$ and
    $X^{\prime\prime}_{k+1}=\min\{X^{\prime\prime}_{k}+\frac{\eta L}{b}\bm{1}_{\{r(k)=l\}}+Z^{\prime}_{k+1},\frac{D}{2}\}\,.$
    Then with the same events $E$ and $E_{0}$, $\mathbb{P}(X_{t}\in E)=1/2$ and
    $\mathbb{P}(X^{\prime}_{k}\leq z)$ $\leq\mathbb{P}(X^{\prime\prime}_{k}\leq z)\text{ for all }z$
    $\mathbb{P}(E_{0}^{c})$ $\leq\exp(-\frac{(0.1D)^{2}}{2\times l\lceil 0.8\frac{Db}{\eta L}\rceil\times(\eta\sigma)^{2}})\leq\exp(-\frac{0.01}{5.6}\mu_{0}^{2})$
    and conditioning on $E_{0}$, $X^{\prime\prime}_{t}=-\tfrac{D}{2}+\tfrac{\eta L}{b}\times\lceil 0.8\tfrac{Db}{\eta L}\rceil\geq 0.3D+\sum_{j=t_{0}}^{t}Z_{j}$; the rest are identical.
Theorem C.17.
In the setting of Theorem 4.5, any valid $f$-DP lower bound for NoisyCGD satisfies
$G(\mu)\geq f$
where $\mu=\frac{L}{b\sigma}\sqrt{\frac{1-c^{lE}}{1+c^{lE}}\frac{1-c^{2}}{(1-c^{l})^{2}}}$.
Proof.
Consider the loss functions in the proof of Theorem 4.2, with $X_{0}=0,\mathcal{K}=\mathbb{R}^{d}$ and $i^{*}\in B_{l}$. By direct calculation $X_{lE}=\mathcal{N}(0,\frac{1-c^{2lE}}{1-c^{2}}\eta^{2}\sigma^{2}I_{d})$ and $X^{\prime}_{lE}=\mathcal{N}(\frac{\eta L}{b}\frac{1-c^{lE}}{1-c^{l}}v,\frac{1-c^{2lE}}{1-c^{2}}\eta^{2}\sigma^{2}I_{d})$, implying $T(X_{lE},X^{\prime}_{lE})=G(\mu)$ with $\mu$ as stated. ∎